# algorithmic

### v1

@@ D.1 Bayes-optimal priors and Gaussian noise channels
The general state evolution depends on the noise channels through $\hat{\Delta}^{z},\tilde{\Delta}^{z}$, and $\bar{R}^{z}$. Using that for the Gaussian channel $P_{\mathrm{out}}(z_{ij},w^{z}_{i},v^{z}_{j})=\mathcal{N}(\lambda_{z}\frac{w^{z}_{i}\;v^{z}_{j}}{\sqrt{n_{z}}},\;\sigma^{2}_{\xi^{z}})$ we have $S^{z}_{ij}=\frac{z_{ij}}{\sigma^{2}_{\xi^{z}}}$ and $R^{z}_{ij}=\frac{z_{ij}^{2}}{\sigma^{4}_{\xi^{z}}}-\frac{1}{\sigma^{2}_{\xi^{z}}}$, we find by plugging into (D.11), (D.16) and (D.19) that
$\hat{\Delta}^{z}$ $=\tilde{\Delta}^{z}=\sigma^{2}_{\xi^{z}}$ (D.34)
$\bar{R}^{z}$ $=0.$ (D.35)
Furthermore, for Bayes-optimal priors we can use the Nishimori identity (23) if the state evolution follows the Nishimori line (Nishimori, 2001). Due to the symmetry spontaneously broken at initialization, as discussed in Section 2.4, $f^{x}_{\mathrm{in}}(K_{i},J_{i})$ is the mean of the local posterior distribution $\mathcal{W}(x,K_{i},J_{i})$ with broken symmetry estimating ±$\pm x^{0}_{i}$, depending on the sign of the node average $\mathbb{E}_{i}[x_{i}^{0}f^{x}_{\mathrm{in}}(K_{i},J_{i})$. Conditioned on the $\pm$ direction of broken symmetry, $f^{x}_{\mathrm{in}}(K,J|\pm)$ has nonzero mean so that self-averaging (D.4) applies and node and disorder average can be exchanged. Then we have the not obvious application
$\mathbb{E}_{x^{0}}[(f^{x}_{\mathrm{in}}(K,J|\pm))^{2}]$ $=\mathbb{E}_{x^{0}}[\mathbb{E}_{\mathcal{W}_{K,J,\pm}}(x)\mathbb{E}_{\mathcal{W}_{K,J,\pm}}(x)]$ (D.36)
$=\mathbb{E}_{x^{0}}[\mathbb{E}_{x_{1},x_{2}\sim\mathcal{W}_{K,J,\pm}}(x_{1}x_{2})]$ (D.37)
$=\mathbb{E}_{x^{0}}[\mathbb{E}_{x_{2}\sim\mathcal{W}_{K,J,\pm}}(\pm x^{0}x_{2})]$ (D.38)
=±$=\pm\mathbb{E}_{x^{0}}[x^{0}\;f^{x}_{\mathrm{in}}(K,J|\pm)]$ (D.39)
$=\left|\mathbb{E}_{x^{0}}[x^{0}\;f^{x}_{\mathrm{in}}(K,J)]\right|,$ (D.40)
where $K,J$ of course depend on $x^{0}$ and in the last step we could exchange the $\pm$ condition for the absolute value. Thus (D.26) and (D.27) yield
$Q^{z,t}_{w}$ $=|M^{z,t}_{w}|$ (D.41)
$Q^{z,t}_{v}$ $=|M^{z,t}_{v}|.$ (D.42)
@@
$\mathbb{E}_{x^{0}}[(f^{x}_{\mathrm{in}}(K,J))^{2}+\partial_{J}f^{x}_{\mathrm{in}}(K,J)]=\mathbb{E}_{x^{0}}[\mathbb{E}_{\mathcal{W}_{K,J}}(xx)]=\mathbb{E}_{x^{0}}[x^{0}x^{0}],$ (D.43)
@@
$Q^{z,t}_{w/v}+\Sigma^{z,t}_{w/v}$ $=\sigma^{2}_{w^{z}/v^{z}}$ (D.44)
The last relation is not needed for the Gaussian channel case, as the terms involving $\Sigma^{z,t}_{w/v}$ vanish anyways due to $\bar{R}^{z}=0$, (D.35). In total, using (D.34),(D.35),D.41, the state evolution simplifies to the form given in Equations 25 and 24.
@@ D.2 Fully Gaussian, and Rademacher-Bernoulli models
With Bayes-optimal, Gaussian priors and Gaussian noise channels, the expectations over both the prior and the source term give a simple closed form. We use the short hands $\tilde{M}^{z,t}_{w/v}=\frac{\lambda^{2}_{z}}{\sigma^{2}_{\xi^{z}}}M^{z,t}_{w/v}$ as introduced also below (24). The denoising function (B.11) being a linear function in $J$, the average over $J^{w,t}_{z}$ in (25) is given by the respective mean $\mathbb{E}_{J^{w,t}_{z}}\left[J^{w,t}_{z}\right]$, leading to
$\mathbb{E}_{J^{w,t}_{z}}\left[f^{w}_{\mathrm{in}}\left(\alpha_{z}|\tilde{M}^{z,t}_{v}|,J^{w,t}_{z}\right)\right]$ $=\frac{\alpha_{z}\tilde{M}^{z,t}_{v}w^{0}_{z}}{\alpha_{z}|\tilde{M}^{z,t}_{v}|+\sigma^{-2}_{w^{z}}}$ (D.45)
@@
$\quad+v_{\bar{z}}^{0}\tilde{M}^{\bar{z},t-1}_{w}\tilde{\Sigma}_{z\bar{z}}\left(|\tilde{M}^{X,t-1}_{w}|,|\tilde{M}^{Y,t-1}_{w}|\right).$ (D.46)
@@
$\quad+c_{v}\tilde{M}^{\bar{z},t-1}_{w}\tilde{\Sigma}_{z\bar{z}}\left(|\tilde{M}^{X,t-1}_{w}|,|\tilde{M}^{Y,t-1}_{w}|\right).$ (D.47)
$M^{z,t}_{w}$ $=\frac{\alpha_{z}\tilde{M}^{z,t}_{v}\sigma^{2}_{w^{z}}}{\alpha_{z}|\tilde{M}^{z,t}_{v}|+\sigma^{-2}_{w^{z}}}$ (D.48)
When changing to a Rademacher-Bernoulli (sparse) prior on $w^{z}$ while $v^{X},v^{Y}$ remains jointly Gaussian, Equation D.47 remains the same. The expectation over the sparse prior in the $M^{z,t}_{w}$ update (25) simply gives a sum of three terms which is omitted here for brevity. Then only the Gaussian integral over the source term $J^{w,t}_{z}$ must be computed numerically.
@@ D.3 Bethe free energy in general case
Based on the form of the SE equations in Appendix D and in analogy to the replica calculation of Lesieur et al. (2017) in their Appendix C, we read off the Bethe free energy, which corresponds to the free energy obtained by a replica-symmetric Ansatz, and which we state here without lengthy derivation
$\Phi^{RS}(\{M,Q,\Sigma\})$ =$=\sum_{\{z\}}\left(\frac{\lambda_{z}\lambda^{0}_{z}}{\hat{\Delta}^{z}}M_{w^{z}}M_{v^{z}}-\frac{\lambda^{2}_{z}}{2\tilde{\Delta}^{z}}Q_{w^{z}}Q_{v^{z}}\right)$
+$\quad+\sum_{\{z\}}\left(\lambda_{z}^{2}\bar{R}^{z}(Q^{z}_{w}+\Sigma^{z}_{w})(Q^{z}_{v}+\Sigma^{z}_{v})\right)$
$\quad-\sum_{\{z\}}\frac{1}{\alpha^{z}}\mathbb{E}_{w_{z},J^{w}_{z}}\left[\log\mathcal{Z}_{w}(K^{w}_{z},J^{w}_{z})\right]$ (D.49)
@@
Here $\mathcal{Z}(K,J)$ are the normalizations of the tilted priors $\mathcal{W}(K,J)$ defined in (10). We can interpret the last two lines of (D.49) as the energetic terms and the first two lines as the additional entropic contributions arising from the introduction of the order parameters after integrating out the Fourier variables. The relation to state evolution is that the stationarity condition
$\vec{\nabla}_{\{M,Q,\Sigma\}}\Phi^{RS}\overset{!}{=}0$ (D.50)
gives back exactly the SE equations (D.24)-(D.33).
@@ D.4 Bethe free energy for Rademacher-Bernoulli prior and Gaussian channels
The Bethe free energy (D.49) simplifies to
$\Phi^{RS}(\{M\})=\sum_{\{z\}}\frac{1}{2}M_{w^{z}}\tilde{M}_{v^{z}}-\sum_{\{z\}}\frac{1}{\alpha^{z}}\overline{\log\mathcal{Z}_{w^{z}}}-\overline{\log\mathcal{Z}_{v}}$ (D.51)
@@
$\overline{\log\mathcal{Z}_{v}}$ $=\mathbb{E}_{(v_{0}^{X},v_{0}^{Y}),J_{X},J_{Y}}\left[\frac{1}{2}\log\det\tilde{\Sigma}_{K}+\frac{1}{2}J^{T}\tilde{\Sigma}_{K}J\right]$ (D.52)
$=\frac{1}{2}\log\det\tilde{\Sigma}_{K}$ (D.53)
$\quad+\frac{1}{2}\sum_{\{z\}}(\tilde{M}^{2}_{w^{z}}\sigma^{2}_{v^{z}}+|\tilde{M}_{w^{z}}|)\tilde{\Sigma}^{zz}_{K}$ (D.54)
$\quad+\frac{1}{2}\tilde{M}_{w^{X}}\tilde{M}_{w^{Y}}c_{v}(\tilde{\Sigma}^{XY}_{K}+\tilde{\Sigma}^{YX}_{K}).$ (D.55)
@@
$\overline{\log\mathcal{Z}_{w^{z}}}$ $=\mathbb{E}_{w^{z},J_{z}}\left[\log(\rho_{z}\cosh(J_{z})e^{-\frac{1}{2}K_{z}}+1-\rho_{z})\right],$ (D.56)
@@ Appendix E Supplementary figures
Figure S2: Phase diagram of weak recovery threshold as in Figure 5, but also showing Bayes-optimal performance specific to the model with Gaussian priors and Gaussian noise channel. The product of cosine similarities, $CS_{w^{z}}CS_{v^{z}}=\frac{M_{w^{z}}M_{v^{z}}}{\sigma_{w^{z}}\sigma_{v^{z}}\sqrt{Q_{w^{z}}Q_{v^{z}}}x}$ as obtained from SE (25,24) is shown. a Both modalities are symmetric, $CS_{w^{z}}CS_{v^{z}}$ is indicated by the green color scale. b The performance achievable in the two modalities differs. $CS_{w^{X}}CS_{v^{X}}$ is shown in blue and $CS_{w^{Y}}CS_{v^{Y}}$ in yellow, mixing to the green color scale on the diagonal which corresponds to the color bar given in panel a. Again the dotted lines indicate the intersection of the two planes.
Figure S3: As Figure 2a, but showing the performance of estimating $v^{z}$ from the same simulations. The $v^{z}$ estimate of linearized AMP in the regime of small noise is not perfect since the operator performs a weighted average of the $v^{X}$ and $v^{Y}$ estimates, as discussed at the end of Section 2.2.

### v2

@@ D.1 Bayes-optimal priors and Gaussian noise channels
The general state evolution depends on the noise channels through $\hat{\Delta}^{z},\tilde{\Delta}^{z}$, and $\bar{R}^{z}$. Using that for the Gaussian channel $P_{\mathrm{out}}(z_{ij},w^{z}_{i},v^{z}_{j})=\mathcal{N}(\lambda_{z}\frac{w^{z}_{i}\;v^{z}_{j}}{\sqrt{n_{z}}},\;\sigma^{2}_{\xi^{z}})$ we have $S^{z}_{ij}=\frac{z_{ij}}{\sigma^{2}_{\xi^{z}}}$ and $R^{z}_{ij}=\frac{z_{ij}^{2}}{\sigma^{4}_{\xi^{z}}}-\frac{1}{\sigma^{2}_{\xi^{z}}}$, we find by plugging into (81), (86) and (89) that
$\hat{\Delta}^{z}$ $=\tilde{\Delta}^{z}=\sigma^{2}_{\xi^{z}}$ (104)
$\bar{R}^{z}$ $=0.$ (105)
Furthermore, for Bayes-optimal priors we can use the Nishimori identity (23) if the state evolution follows the Nishimori line [45]. Due to the symmetry spontaneously broken at initialization, as discussed in Section 2.4, $f^{x}_{\mathrm{in}}(K_{i},J_{i})$ is the mean of the local posterior distribution $\mathcal{W}(x,K_{i},J_{i})$ with broken symmetry estimating $\pm x^{0}_{i}$, depending on the sign of the node average $\mathbb{E}_{i}[x_{i}^{0}f^{x}_{\mathrm{in}}(K_{i},J_{i})$. Conditioned on the $\pm$ direction of broken symmetry, $f^{x}_{\mathrm{in}}(K,J|\pm)$ has nonzero mean so that self-averaging (74) applies and node and disorder average can be exchanged. Then we have the not obvious application
$\mathbb{E}_{x^{0}}[(f^{x}_{\mathrm{in}}(K,J|\pm))^{2}]$ $=\mathbb{E}_{x^{0}}[\mathbb{E}_{\mathcal{W}_{K,J,\pm}}(x)\mathbb{E}_{\mathcal{W}_{K,J,\pm}}(x)]$ (106)
$=\mathbb{E}_{x^{0}}[\mathbb{E}_{x_{1},x_{2}\sim\mathcal{W}_{K,J,\pm}}(x_{1}x_{2})]$ (107)
$=\mathbb{E}_{x^{0}}[\mathbb{E}_{x_{2}\sim\mathcal{W}_{K,J,\pm}}(\pm x^{0}x_{2})]$ (108)
$=\pm\mathbb{E}_{x^{0}}[x^{0}\;f^{x}_{\mathrm{in}}(K,J|\pm)]$ (109)
$=\left|\mathbb{E}_{x^{0}}[x^{0}\;f^{x}_{\mathrm{in}}(K,J)]\right|,$ (110)
where $K,J$ of course depend on $x^{0}$ and in the last step we could exchange the $\pm$ condition for the absolute value. Thus (96) and (97) yield
$Q^{z,t}_{w}$ $=|M^{z,t}_{w}|$ (111)
$Q^{z,t}_{v}$ $=|M^{z,t}_{v}|.$ (112)
@@
$\mathbb{E}_{x^{0}}[(f^{x}_{\mathrm{in}}(K,J))^{2}+\partial_{J}f^{x}_{\mathrm{in}}(K,J)]=\mathbb{E}_{x^{0}}[\mathbb{E}_{\mathcal{W}_{K,J}}(xx)]=\mathbb{E}_{x^{0}}[x^{0}x^{0}],$ (113)
@@
$Q^{z,t}_{w/v}+\Sigma^{z,t}_{w/v}$ $=\sigma^{2}_{w^{z}/v^{z}}$ (114)
The last relation is not needed for the Gaussian channel case, as the terms involving $\Sigma^{z,t}_{w/v}$ vanish anyways due to $\bar{R}^{z}=0$, (105). In total, using (104),(105),111, the state evolution simplifies to the form given in Equations 25 and 24.
@@ D.2 Fully Gaussian, and Rademacher-Bernoulli models
With Bayes-optimal, Gaussian priors and Gaussian noise channels, the expectations over both the prior and the source term give a simple closed form. We use the short hands $\tilde{M}^{z,t}_{w/v}=\frac{\lambda^{2}_{z}}{\sigma^{2}_{\xi^{z}}}M^{z,t}_{w/v}$ as introduced also below (24). The denoising function (59) being a linear function in $J$, the average over $J^{w,t}_{z}$ in (25) is given by the respective mean $\mathbb{E}_{J^{w,t}_{z}}\left[J^{w,t}_{z}\right]$, leading to
$\mathbb{E}_{J^{w,t}_{z}}\left[f^{w}_{\mathrm{in}}\left(\alpha_{z}|\tilde{M}^{z,t}_{v}|,J^{w,t}_{z}\right)\right]$ $=\frac{\alpha_{z}\tilde{M}^{z,t}_{v}w^{0}_{z}}{\alpha_{z}|\tilde{M}^{z,t}_{v}|+\sigma^{-2}_{w^{z}}}$ (115)
@@
$\quad+v_{\bar{z}}^{0}\tilde{M}^{\bar{z},t-1}_{w}\tilde{\Sigma}_{z\bar{z}}\left(|\tilde{M}^{X,t-1}_{w}|,|\tilde{M}^{Y,t-1}_{w}|\right).$ (116)
@@
$\quad+c_{v}\tilde{M}^{\bar{z},t-1}_{w}\tilde{\Sigma}_{z\bar{z}}\left(|\tilde{M}^{X,t-1}_{w}|,|\tilde{M}^{Y,t-1}_{w}|\right).$ (117)
$M^{z,t}_{w}$ $=\frac{\alpha_{z}\tilde{M}^{z,t}_{v}\sigma^{2}_{w^{z}}}{\alpha_{z}|\tilde{M}^{z,t}_{v}|+\sigma^{-2}_{w^{z}}}$ (118)
When changing to a Rademacher-Bernoulli (sparse) prior on $w^{z}$ while $v^{X},v^{Y}$ remains jointly Gaussian, Equation 117 remains the same. The expectation over the sparse prior in the $M^{z,t}_{w}$ update (25) simply gives a sum of three terms which is omitted here for brevity. Then only the Gaussian integral over the source term $J^{w,t}_{z}$ must be computed numerically.
@@ D.3 Bethe free energy in general case
Based on the form of the SE equations in D and in analogy to the replica calculation in [17] Appendix C, we read off the Bethe free energy, which corresponds to the free energy obtained by a replica-symmetric Ansatz, and which we state here without lengthy derivation
$\Phi^{RS}(\{M,Q,\Sigma\})$ $=\sum_{\{z\}}\left(\frac{\lambda_{z}\lambda^{0}_{z}}{\hat{\Delta}^{z}}M_{w^{z}}M_{v^{z}}-\frac{\lambda^{2}_{z}}{2\tilde{\Delta}^{z}}Q_{w^{z}}Q_{v^{z}}\right)$
$\quad+\sum_{\{z\}}\left(\lambda_{z}^{2}\bar{R}^{z}(Q^{z}_{w}+\Sigma^{z}_{w})(Q^{z}_{v}+\Sigma^{z}_{v})\right)$
$\quad-\sum_{\{z\}}\frac{1}{\alpha^{z}}\mathbb{E}_{w_{z},J^{w}_{z}}\left[\log\mathcal{Z}_{w}(K^{w}_{z},J^{w}_{z})\right]$ (119)
@@
Here $\mathcal{Z}(K,J)$ are the normalizations of the tilted priors $\mathcal{W}(K,J)$ defined in (10). We can interpret the last two lines of (119) as the energetic terms and the first two lines as the additional entropic contributions arising from the introduction of the order parameters after integrating out the Fourier variables. The relation to state evolution is that the stationarity condition
$\vec{\nabla}_{\{M,Q,\Sigma\}}\Phi^{RS}\overset{!}{=}0$ (120)
gives back exactly the SE equations (94)-(103). For a rigorous analogue of the replica free energy framework to obtain the posterior mean overlap, we refer the reader to the derivation of mutual information and MMSE in the MTP by Reeves [42].
@@ D.4 Bethe free energy for Rademacher-Bernoulli prior and Gaussian channels
The Bethe free energy (119) simplifies to
$\Phi^{RS}(\{M\})=\sum_{\{z\}}\frac{1}{2}M_{w^{z}}\tilde{M}_{v^{z}}-\sum_{\{z\}}\frac{1}{\alpha^{z}}\overline{\log\mathcal{Z}_{w^{z}}}-\overline{\log\mathcal{Z}_{v}}$ (121)
@@
$\overline{\log\mathcal{Z}_{v}}$ $=\mathbb{E}_{(v_{0}^{X},v_{0}^{Y}),J_{X},J_{Y}}\left[\frac{1}{2}\log\det\tilde{\Sigma}_{K}+\frac{1}{2}J^{T}\tilde{\Sigma}_{K}J\right]$ (122)
$=\frac{1}{2}\log\det\tilde{\Sigma}_{K}$ (123)
$\quad+\frac{1}{2}\sum_{\{z\}}(\tilde{M}^{2}_{w^{z}}\sigma^{2}_{v^{z}}+|\tilde{M}_{w^{z}}|)\tilde{\Sigma}^{zz}_{K}$ (124)
$\quad+\frac{1}{2}\tilde{M}_{w^{X}}\tilde{M}_{w^{Y}}c_{v}(\tilde{\Sigma}^{XY}_{K}+\tilde{\Sigma}^{YX}_{K}).$ (125)
@@
$\overline{\log\mathcal{Z}_{w^{z}}}$ $=\mathbb{E}_{w^{z},J_{z}}\left[\log(\rho_{z}\cosh(J_{z})e^{-\frac{1}{2}K_{z}}+1-\rho_{z})\right],$ (126)
@@ Appendix E Supplementary figures
Figure S2: Phase diagram of algorithmic recovery threshold as in Figure 2, but also showing the Bayes-optimal performance specific to the model with Gaussian priors and Gaussian noise channel. The product of cosine similarities, $CS_{w^{z}}CS_{v^{z}}=\frac{M_{w^{z}}M_{v^{z}}}{\sigma_{w^{z}}\sigma_{v^{z}}\sqrt{Q_{w^{z}}Q_{v^{z}}}x}$ as obtained from SE (25,24) is shown. a Both modalities are symmetric, $CS_{w^{z}}CS_{v^{z}}$ is indicated by the green color scale. b The performance achievable in the two modalities differs. $CS_{w^{X}}CS_{v^{X}}$ is shown in blue and $CS_{w^{Y}}CS_{v^{Y}}$ in yellow, mixing to the green color scale on the diagonal which corresponds to the color bar given in panel a. Again the dotted lines indicate the intersection of the two planes.
Figure S3: As Figure 3, but showing the performance of estimating $v^{z}$ from the same simulations. The $v^{z}$ estimate of linearized AMP in the regime of small noise is not perfect since the operator performs a weighted average of the $v^{X}$ and $v^{Y}$ estimates, as discussed at the end of Section $2.$